OpenStudy (anonymous):

A country's population in 1993 was 94 million. in 1999 in was 99 million. estimate the population in 2005 using the exponential growth formula. round your answer to the nearest millionth. P=Ae^kt

OpenStudy (anonymous):

@myko

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

Nt=N0*e(r*t) Nt = Future Population N0 = Starting population r = growth rate t= number of years. Note: e is a constant equal to 2.71828 Read more: http://wiki.answers.com/Q/What_is_the_exponential_formula_for_population_growth_on_earth#ixzz1zrTB2WH7 now,can u solve for this??

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

P=Ae^kt, this is the formula for population growth in years. There is A and k constants that you need to find first. For that let's use the info we got: 1993 was 94 million. in 1999 in was 99 million. Let's take year 1993 like year 0. So A will be 94. Then in year 1999, it means 6 years later population is 99 mil. 99=94e^6k from here 99/94=e^6k Ln 99/94 = 6k k= 1/6Ln 99/94 Now for year 2005, it means 12 yers after 1993: P = 94 e^(12/6Ln(99/94))=94*(99/94)^2 = 104,265957

OpenStudy (anonymous):

@JGAZAY7